Write a system of two equations in two unknowns. Solve the system by using the substitution method. « Bill and Sue together earn $76,433 per year. If Bill earns $15,000 less than twice Sue’s salary, then how much does each person earn per year? »

Answer by checkley77(12844)

B+S=76,433
B=2S-15,000
(2S-15,000)+S=76,433
2S-15,000+S=76,433
3S=76.433+15.000
3S=91,433
S=91,433/3
S=$30,477.67 IS SUE'S SALARY.
76,433-30,477.67=45,955.33 IS BILL'S SALARY
PROOF:
45,955.33=2*30,477.67-15,000
45,955.33=60,955.33-15,000
45,955.33=45,955.33